Donnell took out a college loan of $3000. He will be charged 8% simple interest on the loan.
A) How much interest will be due on the loan in three months?
B) How much interest will be due on the loan in three years?

Answer by mananth(16949)

Donnell took out a college loan of $3000. He will be charged 8% simple interest on the loan.
A) How much interest will be due on the loan in three months?
B) How much interest will be due on the loan in three years?
..
P= 3000
r=8% = (0.08)
time = 3/12 = 1/4 year
..
I= 3000*0.08*1/4
= $60 after 3 months.
..
per year interest = 60 *4 = $240
for 3years it will be 3 * 240 = $720